# Leveraging LLM Tutoring Systems for Non-Native English Speakers in Introductory CS Courses

### v1

@@ 6. Limitations
Another limitation is in the use of LLMs with datasets that could potentially be considered too large—where some natural languages may benefit from the most advanced techniques when others may not (Bender et al., 2021). Furthermore, LLMs have been found to exhibit stereotypes and cultural biases towards certain groups of people outside of Anglophonic settings (Maina et al., 2024). Future work should actively test the LLMs used in their study on whether it displays biases for particular natural languages or cultural biases and clearly state those biases.
Additionally, we acknowledge that the introduction of CodeHelp was through a class forum post. This may contribute to the system’s visibility and usage issues found in Section 4.2.1. We suggest future implementations of LLM tutors provide in-class demonstrations of the system highlighting how to use the tool as well as its multilingual capabilities—possibly increasing overall usage and multilingual queries. Alongside the demonstration, an onboarding video and/or text manual for the system may serve as a valuable, asynchronous resource for students with usage issues.
7. Conclusion

### v2

@@ VI Limitations
Another limitation is in the use of LLMs with datasets that could potentially be considered too large—where some natural languages may benefit from the most advanced techniques when others may not [8]. Furthermore, LLMs have been found to exhibit stereotypes and cultural biases towards certain groups of people outside of Anglophonic settings [30]. Future work should actively test the LLMs used in their study on whether it displays biases for particular natural languages or cultural biases and clearly state those biases.
Additionally, we acknowledge that the introduction of CodeHelp was through a class forum post. This may contribute to the system’s visibility and usage issues found in Section IV-B1. We suggest future implementations of LLM tutors provide in-class demonstrations of the system highlighting how to use the tool as well as its multilingual capabilities—possibly increasing overall usage and multilingual queries. Alongside the demonstration, an onboarding video and/or text manual for the system may serve as a valuable, asynchronous resource for students with usage issues.
VII Conclusion
@@
VIII Acknowledgements
This work was supported in part by the NSF Grant #2417374, UCSD Sloan Scholars Fellowship, and Gates Millennium Scholarship. Thanks to Jackelyne García-Villegas for their feedback on this paper.